The pH of 0.50 M HClO is 3.91. Calculate the change in pH when 1.29 g of NaClO is added to 18 mL of 0.50 M HClO. Ignore any changes in volume. The Ka value for HClO is 3.0 x 10-8.
Answer #1

pH = pKa + log (salt) / (acid)

number of moles of NaClO = 1.29 / 74.44 = 0.01733 moles

concentration of NaClO = (1.29 / 74.44) * (1000/18) = 0.9627 mole / L

pH = pKa + log (salt) / (acid)

pH = -log(3 * 10^-8) + log (0.9627 / 0.5 )

pH = 7.807

chane in pH = 7.807 - 3.91 = 3.897